# An Unforgeable Publicly Verifiable Watermark for Large Language Models

### v4

@@ Abstract
Recently, text watermarking algorithms for large language models (LLMs) have been proposed to mitigate the potential harms of text generated by LLMs, including fake news and copyright issues. However, current watermark detection algorithms require the secret key used in the watermark generation process, making them susceptible to security breaches and counterfeiting during public detection. To address this limitation, we propose an unforgeable publicly verifiable watermark algorithm that uses two different neural networks for watermark generation and detection, instead of using the same key at both stages. Meanwhile, the token embedding parameters are shared between the generation and detection networks, which makes the detection network achieve a high accuracy very efficiently. Experiments demonstrate that our algorithm attains high detection accuracy and computational efficiency through neural networks with a minimized number of parameters. Subsequent analysis confirms the high complexity involved in forging the watermark from the detection network. Our code and data are available at https://github.com/THU-BPM/unforgeable_watermark.
@@ 1 Introduction
With the development of current large language models (LLMs), many LLMs, like GPT-4 (OpenAI, 2023) and Claude [^1] https://claude.ai/chat, could rapidly generate human-like texts. This has led to numerous risks, such as the generation of a vast amount of false information on the Internet (Pan et al., 2023), and the infringement of copyrights of creative works (Chen et al., 2023). Therefore, texts generated by LLMs need to be detected and tagged.
@@
In this work, we propose the first unforgeable publicly verifiable watermarking algorithm for large language models (LLMs). Our approach adopts the commonly used watermark schema, which embeds small watermark signals into LLM’s logits during generation Kirchenbauer et al. (2023); Zhao et al. (2023). The main difference is that our algorithms could detect the watermark without the generation key. To conceal the details of the watermark in the detection process, we propose using two separate neural networks for watermark generation and detection, rather than relying on the same secret key for both stages. The unforgeability of our algorithm stems from a computational asymmetry: constructing a watermark generation network from a watermark detection network is notably more complex than the reverse process (Section 4.5). To enhance the accuracy of the detection network while alleviating the complexity of the training process, we have shared the token embedding parameters between the detector and the generator. This provides some prior information to the detector while maintaining unforgeable. Specifically, our watermark generator takes a window of $w$ tokens as input and predicts the watermark signals of the last token. The text detector directly takes all tokens from the text as input and outputs whether the complete text contains the watermark.
@@
  • Introduction of a novel unforgeable publicly verifiable algorithm that employs separate neural networks for watermark generation and detection.
@@ 2 Related work
As large language models (LLMs) generate increasingly high-quality text, detecting and tagging machine-generated content grows more vital. Currently, there are two primary methods for identifying LLM-produced text: text watermarking and classifier-based detection. With watermarking, implicit features are incorporated into the text during generation, allowing specially designed techniques to identify marked passages. Classifier-based detection leaves text generation unchanged and uses classifiers trained to differentiate machine-from human-authored content.
@@
Recent work has explored classifier-based approaches for detecting machine-generated text. These methods typically fine-tune large pre-trained language models like GPT-2 (Radford et al., 2019), BART(Lewis et al., 2019), and RoBERTa (Liu et al., 1907) for binary text classification (Zhan et al., 2023; Mireshghallah et al., 2023). To improve model robustness, some methods incorporate language model features (Su et al., 2023) or adversarial training (Hu et al., 2023). However, the fundamental detectability of machine-generated text remains an open question. Some argue that with sufficient data collection, high-performing detectors are achievable (Chakraborty et al., 2023). Others contend that as language models become more advanced, the performance gap between detectors and random classifiers may diminish (Sadasivan et al., 2023).
@@
Compared to the classifier-based methods, text watermarking is more explainable due to that the implicit features are already injected in the text. There are typically two categories of text watermarking methods. The first is to add a watermark to the existing text. For example, Abdelnabi & Fritz (2021) designed a data-hiding network to embed watermark information in the text, and utilized a data-revealing network to recover the embedded information. Yoo et al. (2023) injected the watermark by substituting some words in the text. However, adding a watermark to the existing text struggles to keep the semantics of the text unchanged which limits its use in real-world scenarios. Another line of methods is injecting the watermark during the text decoding process. Christ et al. (2023) used a pre-selected value sequence to sample the tokens and subsequently detected the watermark by observing the correlation between the preset pseudorandom numbers and the generated tokens. The work of Kuditipudi et al. (2023) introduced Levenshtein distance to measure the distance between generated text and random number sequences, improving the robustness of watermarks. Kirchenbauer et al. (2023) divided the vocabulary into red and green lists and preferred to generate tokens from the green list. Zhao et al. (2023) enhanced the robustness of this approach by using a global fixed red-green vocabulary. Lee et al. (2023) designed a watermarking method for low-entropy code generation scenarios. However, the above methods require the key from the watermark generation phase during the detection, which exposes the watermark’s key when made accessible for public detection, consequently making it susceptible to forgery. In this work, we propose the first unforgeable publicly verifiable watermark to alleviate these issues.
@@ 3 Problem definition
Watermarking Algorithm: This construct comprises two synergistic algorithms—the watermark generation algorithm and the watermark detection algorithm.
  • Watermark Generation: This can be seen as a subtle modification to a language model’s probability distribution. If $\hat{\mathcal{M}}$ denotes a watermarked language model, the token prediction probability is expressed as: $\hat{P}_{n}:=\hat{\mathcal{M}}(\bm{x}_{0:n-1})$.
  • Watermark Detection: This algorithm takes a text sequence $\bm{x}=[x_{0},\dots,x_{n}]$ and determines the presence of a watermark. There’s a one-to-one correspondence between the watermark detection model, denoted as $\mathcal{D}$, and the watermarked language model $\hat{\mathcal{M}}$.

### v7

@@ Abstract
Recently, text watermarking algorithms for large language models (LLMs) have been proposed to mitigate the potential harms of text generated by LLMs, including fake news and copyright issues. However, current watermark detection algorithms require the secret key used in the watermark generation process, making them susceptible to security breaches and counterfeiting during public detection. To address this limitation, we propose an unforgeable publicly verifiable watermark algorithm named UPV that uses two different neural networks for watermark generation and detection, instead of using the same key at both stages. Meanwhile, the token embedding parameters are shared between the generation and detection networks, which makes the detection network achieve a high accuracy very efficiently. Experiments demonstrate that our algorithm attains high detection accuracy and computational efficiency through neural networks. Subsequent analysis confirms the high complexity involved in forging the watermark from the detection network. Our code is available at https://github.com/THU-BPM/unforgeable_watermark. Additionally, our algorithm could also be accessed through MarkLLM (Pan et al., 2024)  [^1] https://github.com/THU-BPM/MarkLLM.
@@ 1 Introduction
With the development of current large language models (LLMs), many LLMs, like GPT-4 (OpenAI, 2023) and Claude, could rapidly generate human-like texts. This has led to numerous risks, such as the generation of a vast amount of false information on the Internet (Pan et al., 2023), and the infringement of copyrights of creative works (Chen et al., 2023). Therefore, texts generated by LLMs need to be detected and tagged.
@@
In this work, we propose UPV, the first unforgeable publicly verifiable watermarking algorithm for LLMs. Our approach adopts the commonly used watermark schema, which embeds small watermark signals into LLM’s logits during generation Kirchenbauer et al. (2023); Zhao et al. (2023). The main difference is that our algorithms could detect the watermark without the generation key. To conceal the details of the watermark in the detection process, we propose using two separate neural networks for watermark generation and detection, rather than relying on the same secret key for both stages. The unforgeability of our algorithm stems from a computational asymmetry: constructing a watermark generation network from a watermark detection network is notably more complex than the reverse process (Section 4.5). To enhance the accuracy of the detection network while alleviating the complexity of the training process, we have shared the token embedding parameters between the detector and the generator. This provides some prior information to the detector while maintaining unforgeable. Specifically, our watermark generator takes a window of $w$ tokens as input and predicts the watermark signals of the last token. The text detector directly takes all tokens from the text as input and outputs whether the complete text contains the watermark.
@@
  • Introduction of a novel unforgeable publicly verifiable algorithm named UPV that employs separate neural networks for watermark generation and detection.
@@ 2 Related work
LLMs have been successfully applied to many tasks (Liu et al., 2023a; Hu et al., 2023b). However, the potential for misuse of LLMs still remains. To mitigate these misuses, strategies such as red teaming (Liu et al., 2022) and safety alignment (Ji et al., 2023) can be employed. Nonetheless, not all forms of misuse can be addressed through these methods, making it imperative to detect and track text generated by LLMs. Currently, there are two primary methods for identifying LLM-produced text: text watermarking and classifier-based detection.
@@
Classifier-based approaches typically fine-tune LLMs like GPT-2 (Radford et al., 2019), BART(Lewis et al., 2019), and RoBERTa (Liu et al., 1907) for binary text classification (Zhan et al., 2023; Mireshghallah et al., 2023). Some methods incorporate language model features (Su et al., 2023) or adversarial training (Hu et al., 2023a). However, the fundamental detectability of machine-generated text remains an open question. Some argue that with sufficient data collection, high-performing detectors are achievable (Chakraborty et al., 2023). Sadasivan et al. (2023) contend that as LLMs become more advanced, the performance gap between detectors and random classifiers may diminish.
@@
Compared to the classifier-based methods, text watermarking (Liu et al., 2023c) is more explainable. There are typically two categories of text watermarking methods. The first is to add a watermark to the existing text. For example, Abdelnabi & Fritz (2021) designed a data-hiding network to embed watermark information in the text, and utilized a data-revealing network to recover the embedded information. Yoo et al. (2023) injected the watermark by substituting some words in the text. However, adding a watermark to the existing text struggles to keep the semantics of the text unchanged which limits its use in real-world scenarios. Another line of methods is injecting the watermark during the text decoding process. Christ et al. (2023) used a pre-selected value sequence to sample the tokens and subsequently detected the watermark by observing the correlation between the preset pseudorandom numbers and the generated tokens. The work of Kuditipudi et al. (2023) introduced Levenshtein distance to measure the distance between generated text and random number sequences, improving the robustness of watermarks. Kirchenbauer et al. (2023) divided the vocabulary into red and green lists and preferred to generate tokens from the green list. Zhao et al. (2023) enhanced the robustness of this approach by using a global fixed red-green vocabulary. Liu et al. (2023b) achieve a balance between robustness and security through semantic invariant hashing. Lee et al. (2023) designed a watermarking method for low-entropy code generation scenarios. However, the above methods require the key from the watermark generation phase during the detection, which exposes the watermark’s key when made accessible for public detection. In this work, we propose the first unforgeable publicly verifiable watermark to alleviate these issues.
@@ 3 Problem definition
Watermarking Algorithm comprises watermark generation and watermark detection. Watermark Generation: This can be seen as a subtle modification to a language model’s probability distribution. If $\hat{\mathcal{M}}$ denotes a watermarked language model, the token prediction probability is expressed as: $\hat{P}_{n}:=\hat{\mathcal{M}}(\bm{x}_{0:n-1})$. Watermark Detection: This algorithm takes a text sequence $\bm{x}=[x_{0},\dots,x_{n}]$ and determines the presence of a watermark. There’s a one-to-one correspondence between the watermark detection model, denoted as $\mathcal{D}$, and the watermarked language model $\hat{\mathcal{M}}$.
@@ 6 Conclusion
7 Acknowledgments
This work is supported by the National Nature Science Foundation of China (No. 62021002), Tsinghua BNRist, and the Beijing Key Laboratory of Industrial Bigdata System and Application. Additionally, it receives support from the Beijing Natural Science Foundation under grant number QY23117. Moreover, the work is partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 14222922, RGC GRF No. 2151185). This work is also supported in part by NSF under grant III-2106758.
At the same time, we would like to express our gratitude to the anonymous ICLR reviewers PMgN, CjAu, seaW, and kypk, as well as the Area Chair vQPR, for their valuable feedback, which played a significant role in improving the quality of our paper. We especially appreciate reviewer CjAu, who helped us clarify the definition of our methods and provided many valuable suggestions for revisions, greatly enhancing the quality of our paper.